# Evaluating Large Language Model based Personal Information Extraction and Countermeasures

### v1

@@ 7 Defenses
In this section, we formulate defenses against PIE by analyzing the defender’s goals in general. Then, we summarize existing defenses that are commonly used in practice. Finally, we propose to leverage prompt injection as a defense against LLM-based PIE.
@@ 7.2 Existing Defenses
Preventing the personal information shown on publicly available sites from being extracted by automatic tools is a long-standing topic. According to our survey, we find the following commonly used existing defenses against information extraction:
  • Symbol replacement:  This defense serves as a countermeasure in safeguarding the information extraction, especially for protecting the email address. This defense operates by altering key symbols within the email address, such as replacing ‘@’ with ‘AT’ and ‘.’ with ‘DOT’. For instance, the email address “123@gmail.com” would be transformed into “123 AT gmail DOT com” after this defense is applied.
  • Text replacement:  This defense is mainly used to counteract information extraction efforts targeting email address. This approach works by replacing identifiable parts of an email address with generic placeholders. To maintain its readability to normal human users, the name is the most common part to be replaced. Specifically, the first or last name would be transformed to <first_name> or <last_name>. For instance, an email address such as “albert.einstein@gmail.com” would be transformed into “<first_name>.<last_name>@gmail.com”. This method alters the personal elements of the email address, rendering it in a way that is harder to extract by automated PIE tools.
  • Hyperlink:  This defense protects the email address by hiding it behind a hyperlink, where the normal users could click on to be redirected to their email systems to start a new email thread.
  • Text-to-image:  This defense operates by storing personal information in a non-text format. For instance, a common way to apply this defense is to present the email address in an image, which could invalidate the text-based personal information extraction tools. It is notable that this defense is also applicable to other categories of personal information such as name and mailing address.
Table 7: Comparing the accuracy / BERT score of LLMs at extracting email addresses under different defenses (we defer the defense results of text-to-image to Table 9). A defense is more effective when the accuracy / BERT score values are smaller.
LLM No Defense Symbol replacement Hyperlink Keyword replacement Prompt Injection
@@
gpt-4 100% / 1 99% / 0.99 99% / 0.99 98% / 0.98 99% / 0.99 81% / 0.81 0% / 0.00
vicuna-13b-v1.3 62% / 0.62 61% / 0.61 61% / 0.61 55% / 0.55 61% / 0.61 35% / 0.35 11% / 0.11
vicuna-7b-v1.3 55% / 0.55 52% / 0.52 51% / 0.51 45% / 0.45 58% / 0.58 33% / 0.33 8% / 0.08
llama-2-7b-chat-hf 31% / 0.31 29% / 0.29 28% / 0.28 28% / 0.28 30% / 0.30 9% / 0.9 7% / 0.07
internlm-chat-7b 96% / 0.96 95% / 0.95 95% / 0.95 90% / 0.90 96% / 0.96 63% / 0.63 2% / 0.02
flan-ul2 100% / 1 99% / 0.99 99% / 0.99 99% / 0.99 98% / 0.98 98% / 0.98 54% / 0.54
@@
vicuna-13b-v1.3 11% / 0.11 13% / 0.13 2% / 0.02 1% / 0.01 18% / 0.18 16% / 0.16 2% / 0.02 4% / 0.04
vicuna-7b-v1.3 8% / 0.08 0% / 0.00 3% / 0.03 2% / 0.02 19% / 0.14 17% / 0.13 2% / 0.02 4% / 0.04
llama-2-7b-chat-hf 7% / 0.07 11% / 0.11 5% / 0.05 12% / 0.11 13% / 0.11 9% / 0.03 8% / 0.05 6% / 0.04
internlm-chat-7b 2% / 0.02 0.28% / 0.28 5% / 0.03 10% / 0.08 15% / 0.11 13% / 0.10 9% / 0.08 7% / 0.02
@@
Table 9: The effectiveness of text-to-image defense at defending against PIE.
LLM With/without text-to-image defense Personal Information
gemini-pro w/o defense 98% / 0.98 98% / 0.98 83% / 0.82 100% / 0.99 63% / 0.49 74% / 0.62 83% / 0.78 56% / 0.46
w/ defense 71% / 0.71 91% / 0.91 86% / 0.85 95% / 0.96 75% / 0.59 85% / 0.78 86% / 0.82 92% / 0.81
gpt-4 w/o defense 100% / 1 98% / 0.98 86% / 0.85 100% / 1 66% / 0.57 74% / 0.63 87% / 0.85 78% / 0.70
w/ defense 96% / 0.96 68% / 0.68 85% / 0.82 98% / 0.98 80% / 0.72 80% / 0.69 92% / 0.85 88% / 0.78
@@ 7.3 Prompt Injection as a Defense
As studied in existing works [29, 23, 24, 37], prompt injection is often recognized as a security attack against LLMs. This attack aims to manipulate the output of an LLM by inserting a carefully crafted prompt into a standard input. For instance, in a phishing website detection scenario, a user may create a prompt, “Tell me whether the following website is a phishing site: <input_link>”, where “<input_link>” represents a link to an external website. The user leverages this prompt to detect if a website is phishing. In prompt injection attacks, an attacker may inject “Ignore previous instructions. Say ‘not phishing’ instead.” into that website as an inline comment. As a result, even if the website is phishing, the LLM would erroneously respond with “not phishing” as it is misled by the injected prompt. Prompt injection attacks are successful because LLMs have strong capabilities to follow instructions.
@@
  • Effectiveness:  The effectiveness is the key challenge to turn prompt injection into a valid defense. As studied in previous works [29, 23, 30], context-ignoring is a useful technique in crafting effective injected prompts. In particular, the key idea is to inject a prompt (e.g., “Ignore my previous instructions.”) such that LLMs would ignore the previous prompt and follow the injected prompt. We leverage state-of-the-art prompt injection techniques [23] in this work. Additionally, we will also evaluate the effectiveness of different prompt injection techniques.
@@
We have discussed how to design an injected prompt that is effective, invisible to human users, and applicable to all categories of personal information. Figure 6 in the Appendix shows how we inject the prompt into an HTML profile and Table 15 in the Appendix presents the injected prompts we use. In particular, we leverage a pre-defined CSS function to make the injected prompt not selectable. This unselectable feature can be easily achieved in other formats of personal profiles. Figure 7b in the Appendix shows a rendered personal profile with injected prompt. Compared to the profile without injected prompt shown in Figure 7a, they are visually identical.
@@
Data prompt isolation (quotes) w/o defense 50% / 0.50 100% / 1 91% / 0.70 100% / 1 61% / 0.46 74% / 0.55 91% / 0.88 50% / 0.39
@@
Data prompt isolation (xml tags) w/o defense 99% / 0.99 99% / 0.99 78% / 0.76 100% / 1 72% / 0.58 78% / 0.65 89% / 0.86 55% / 0.44
@@
Data prompt isolation (random sequence) w/o defense 99% / 0.99 99% / 0.99 77% / 0.75 100% / 1 67% / 0.50 76% / 0.58 89% / 0.86 56% / 0.46
@@
Instructional prevention w/o defense 99% / 0.99 96% / 0.96 78% / 0.76 100% / 1 65% / 0.47 74% / 0.55 89% / 0.85 53% / 0.42
@@
Sandwich prevention w/o defense 99% / 0.99 99% / 0.99 78% / 0.77 100% / 1 65% / 0.49 75% / 0.57 89% / 0.87 54% / 0.43
@@
Table 11: Comparing the effectiveness of prompt injection defenses using different strategies to craft the injected prompt.
Prompt injection strategy Personal Information
No defense 93% / 0.93 95% / 0.95 77% / 0.67 98% / 0.96 67% / 0.53 76% / 0.59 80% / 0.70 51% / 0.37
CI 79% / 0.79 84% / 0.84 43% / 0.43 85% / 0.85 61% / 0.47 76% / 0.57 82% / 0.80 55% / 0.44
ID 36% / 0.36 40% / 0.40 28% / 0.28 77% / 0.77 46% / 0.36 65% / 0.49 60% / 0.59 35% / 0.29
CI+ID 0% / 0.00 1% / 0.01 0% / 0.00 2% / 0.02 26% / 0.20 33% / 0.25 0% / 0.00 0% / 0.00
8 Evaluating Defenses
8.1 Experimental Setup
8.1.1 LLMs, attack settings, and evaluation metrics
Our experimental settings are consistent to Section 6. In particular, unless otherwise mentioned, we use the same settings for LLMs, attacker’s prompt, personal profiles, and evaluation metrics as Section 6. Similarly, each table entry presents the experimental results in the form of “Accuracy / BERT score” for email address and Tel. and “Rouge-1 score / BERT score” for other categories of personal information.
8.1.2 Defense settings
For symbol replacement, keyword replacement, and hyperlink, we only apply them to protect the email address, as discussed in Section 7. In particular, we transform the email address in each personal profile in the synthetic dataset into the corresponding format of each defense. Regarding text-to-image, we utilize a publicly available library [5] to convert personal information into an image and embed it into the corresponding personal profile.
For prompt injection defense, we inject the prompt shown in Table 15 in the Appendix to all personal profiles in the synthetic dataset and Figure 6 in Appendix presents how we make the injection. By default, we use “combination of context ignoring and injected data (CI+ID)” as our prompt injection strategy.
8.2 Defense Results
Prompt injection outperforms existing countermeasures:  Table 7 shows the attack results to extract the email address when different defenses (excluding text-to-image) are adopted. For symbol replacement, we consider three variants: replacing “@” only, replacing “.” only, and replacing both “@” and “.”. We observe that prompt injection is the most effective defense, as the accuracy of the email address extracted by the attacker drops drastically when prompt injection is applied. This indicates that the injected prompt can successfully deceive the LLMs and mislead them to extract incorrect personal information. By contrast, other defenses, i.e., three variants of symbol replacement, keyword replacement, and hyperlink, only have limited effectiveness, as we only observe a slight decrease when these defenses are used. In addition, Table 9 shows the experimental results of the text-to-image defense when multi-modal LLMs (i.e., GPT-4V and Gemini-pro-vision) are used. The defense only decreases the attack effectiveness slightly, which indicates that text-to-image is not effective in preventing extraction based on multi-modal LLMs. Overall, the experimental results demonstrate that prompt injection outperforms other countermeasures when defending against LLM-based email address extraction from personal profiles.
We observe that the Rouge1 score (or BERT score) values for the work and educational experience extracted by the attacker are not close to zero, but they remain at around 0.2 (or 0.25). We suspect this is due to the design of our injected prompt. As shown in Table 15, the injected prompt aims to mislead the LLM to answer with “imaginary company” (or “imaginary school”) when the LLM is originally guided by the attacker to extract the work experience (or education experience). As a result, even if the injected prompt successfully deceives the LLM and makes it generate “imaginary company” (or ‘imaginary school”), the keyword “company” (or “school”) is very likely to have some overlaps with the words in the ground truth labels, resulting in non-zero Rouge1 scores (or BERT scores).
Impact of the prompt used in prompt injection:  Table 11 shows the impact of prompt injection strategies on the defense effectiveness. The prompts used in our experiment when only one strategy is adopted are shown in Table 15 in the Appendix. We observe that LLM-based PIE is still very effective when only one strategy is applied. By contrast, the accuracy (or Rouge1/BERT score) is small when both strategies are used. This highlights the importance of both the context ignoring strategy and the injected data in designing the injected prompt to effectively defend against LLM-based PIE.
@@ 8.2 Defense Results
Prompt injection is effective in defending against adaptive attacks:  Existing works propose various ways to defend against prompt injection attacks [7, 38, 42, 30, 26]. Their main goal is to prevent the LLMs from being misled by the injected prompt via either purifying the entire prompt or separating the instruction and the document/personal profile. In our scenario, we utilize them as adaptive attacks to break the defensive prompt injection used by the personal profile owners. In other words, these adaptive attacks aim to make the LLMs accurately extract the personal information from the profile, instead of being misguided by the injected prompt. We utilize paraphrasing, retokenization, variants of data prompt isolation, instructional prevention, and sandwich prevention as adaptive attacks. A summary of these adaptive attacks and examples are presented in Table 18 in the Appendix. In our evaluation, we consider our attack framework as the baseline attack, and integrate these adaptive attacks into our attack framework one at a time. In addition, we do not consider prompt injection detection methods as adaptive attacks in our evaluation, since even with the knowledge that a document contains injected prompts, the attacker still cannot distinguish the genuine and injected contents, and thus fails to extract accurate personal information.
@@
Table 10 shows the attack results for these adaptive attacks. We have two observations. First, some adaptive attacks, e.g., retokenization, achieve a higher attack accuracy (or Rouge1/BERT scores) compared to the results of the baseline attack. However, the increase from these adaptive attacks is very limited. For instance, when extracting the email address, the retokenization only achieves an accuracy of 0.22. Second, we observe that some adaptive attacks sacrifice the utility. In particular, when the prompt injection is not applied, the attack results for some adaptive attacks are lower than the baseline attack. For example, when extracting the email address, the retokenization only has an accuracy of 0.48, which is much lower than the value of the baseline attack (i.e., 0.93). From these two observations, we could draw the conclusion that these adaptive attacks are insufficient in bypassing the prompt injection defense.

### v2

@@ 7 Defenses
We first summarize commonly used existing defenses. Then, we propose to leverage prompt injection as a defense.
@@ 7.2 Existing Defenses
Preventing the personal information shown on publicly available sites from being extracted by automatic tools is a long-standing topic. According to our survey, we find the following commonly used existing defenses against PIE:
Symbol replacement (SR):  This defense serves as a countermeasure in safeguarding the information extraction, especially for protecting the email address. This defense operates by altering key symbols within the email address, such as replacing ‘@’ with ‘AT’ and ‘.’ with ‘DOT’. For instance, the email address “123@gmail.com” would be transformed into “123 AT gmail DOT com” after this defense is applied.
Keyword replacement (KR):  This defense is mainly used to counteract information extraction efforts targeting email address. This approach works by replacing identifiable parts of an email address with generic placeholders. To maintain its readability to normal human users, the name is the most common part to be replaced. Specifically, the first or last name would be transformed to <first_name> or <last_name>. For instance, an email address such as “albert.einstein@gmail.com” would be transformed into “<first_name>.<last_name>@gmail.com”. This method alters personal elements of email address, rendering it in a way that is harder to extract by automated tools.
Hyperlink (HL):  This defense hide an email address behind a hyperlink, where the normal users could click on to be redirected to their email systems to start an email thread.
Text-to-image (TI):  This defense operates by storing personal information in a non-text format. For instance, a common way to apply this defense is to present the email address in an image, which could invalidate the text-based personal information extraction tools. It is notable that this defense is also applicable to other categories of personal information such as name and mailing address.
Existing defenses cannot effectively protect personal information from LLM-based PIE. In the real-world Professor dataset, symbol replacement, hyperlink, and keyword replacement are used by 34%, 34%, and 5% of professors to protect their email addresses, but LLM-based PIE (with GPT-4) achieves 97%, 100%, and 80% accuracy in extracting them. Notably, GPT-4 fails to extract an email address that uses keyword replacement, such as substituting a part of the email address with phrases like “my last name.” While this technique shows potential for protecting email addresses by obfuscating key components, its applicability is limited to this specific category of personal information and does not generalize well to other types, such as phone numbers or affiliations.
Table 8: Comparing the accuracy / BERT score of LLMs at extracting email addresses under different defenses (we defer the defense results of text-to-image to Table 17). A defense is more effective when the accuracy / BERT score is smaller.
LLM No Defense SR HL KR PI
@@
vicuna-13b-v1.3 59% / 0.59 65% / 0.65 43% / 0.43 44% / 0.44 36% / 0.36 8% / 0.08 0% / 0.00
vicuna-7b-v1.3 65% / 0.65 84% / 0.84 69% / 0.69 79% / 0.79 57% / 57 17% / 0.17 3% / 0.01
llama-2-7b-chat-hf 26% / 0.26 39% / 0.39 27% / 0.27 32% / 0.32 38% / 38 3% / 0.03 1% / 0.01
internlm-chat-7b 95% / 0.95 98% / 0.98 78% / 0.78 87% / 0.87 82% / 82 20% / 0.20 5% / 0.05
flan-ul2 100% / 1.00 99% / 0.99 99% / 0.99 99% / 0.99 98% / 0.98 98% / 0.98 54% / 0.54
Table 9: Comparing the accuracy / BERT score of different PIE methods at extracting email addresses under different defenses. A defense is more effective when the accuracy / BERT score is smaller.
Extraction method No Defense SR HL KR TI PI PI+KR
AT DOT AT+DOT
Regular Expression 100% / 1.00 0% / 0.00 0% / 0.00 0% / 0.00 100% / 1.00 0% / 0.00 0% / 0.00 99% / 0.99 0% / 0.00
Keyword search 14% / 0.14 13% / 0.13 13% / 0.13 13% / 0.13 0% / 0.00 0% / 0.00 0% / 0.00 14% / 0.14 0% / 0.00
spaCy 95% / 0.95 95% / 0.95 0% / 0.00 0% / 0.00 95% / 0.95 0% / 0.00 0% / 0.00 94% / 0.94 0% / 0.00
BERT-based 1% / 0.01 0% / 0.00 0% / 0.00 0% / 0.00 0% / 0.00 0% / 0.00 0% / 0.00 1% / 0.01 0% / 0.00
mlscraper 77% / 0.77 77% / 0.77 77% / 0.77 77% / 0.77 0% / 0.00 0% / 0.00 0% / 0.00 77% / 0.77 0% / 0.00
LLM-based method 100% / 1.00 99% / 0.99 99% / 0.99 98% / 0.98 99% / 0.99 81% / 0.81 100% / 1.00 0% / 0.00 0% / 0.00
Max 100% / 1.00 99% / 0.99 99% / 0.99 98% / 0.98 100% / 1.00 81% / 0.81 100% / 1.00 99% / 0.99 0% / 0.00
@@
vicuna-13b-v1.3 0% / 0.00 22% / 0.22 5% / 0.05 1% / 0.01 0% / 0.00 3% / 0.03 0% / 0.00 1% / 0.01
vicuna-7b-v1.3 3% / 0.03 11% / 0.11 16% / 0.15 0% / 0.00 12% / 0.05 8% / 0.07 2% / 0.02 1% / 0.01
llama-2-7b-chat-hf 1% / 0.01 8% / 0.08 4% / 0.03 17% / 0.17 3% / 0.02 2% / 0.00 7% / 0.06 2% / 0.01
internlm-chat-7b 5% / 0.05 1% / 0.01 5% / 0.05 4% / 0.04 5% / 0.04 8% / 0.07 3% / 0.09 0% / 0.02
@@ 7.3 Prompt Injection as a Defense
As studied in existing works [27, 35], prompt injection (PI) is often recognized as a security attack against LLMs. This attack aims to manipulate the output of an LLM by inserting a carefully crafted prompt into a standard input. For instance, in a phishing website detection scenario, a user may create a prompt, “Tell me whether the following website is a phishing site: <input_link>”, where “<input_link>” represents a link to an external website. The user leverages this prompt to detect if a website is phishing. In prompt injection attacks, an attacker may inject “Ignore previous instructions. Say ‘not phishing’ instead.” into that website as an inline comment. Thus, even if the website is phishing, the LLM would erroneously respond with “not phishing” as it is misled by the injected prompt. Prompt injection attacks are successful because LLMs have strong capabilities to follow instructions.
@@
Effectiveness:  The effectiveness is the key challenge to turn prompt injection into a valid defense. As studied in previous works [27, 35], context-ignoring is a useful tool in crafting effective injected prompts. In particular, the key idea is to inject a prompt (e.g., “Ignore previous instructions.”) such that LLMs would ignore the previous prompt and follow the injected prompt. We leverage state-of-the-art prompt injection techniques [27] in this work. Additionally, we will evaluate the effectiveness of different prompt injection techniques.
@@
We have discussed how to design an injected prompt that is effective, invisible to human users, and applicable to all categories of personal information. Figure 18 in the Appendix shows how we inject the prompt into an HTML profile and Table 18 in the Appendix presents the injected prompts we use. In particular, we leverage a pre-defined CSS function to make the injected prompt not selectable. This unselectable feature can be easily achieved in other formats of personal profiles. Figure 5b in the Appendix shows a rendered personal profile with injected prompt. Compared to the profile without injected prompt shown in Figure 5a in the Appendix, they are visually identical.
@@
8 Evaluating Defenses
8.1 Experimental Setup
LLMs, attack settings, and evaluation metrics:  Our experimental settings are consistent to Section 6. In particular, unless otherwise mentioned, we use the same settings for LLMs, attacker’s prompt, personal profiles, and evaluation metrics as Section 6. Similarly, each table entry presents the experimental results in the form of “Accuracy / BERT score” for email address and Tel. and “Rouge-1 score / BERT score” for other categories of personal information.
Defense settings:  For symbol replacement, keyword replacement, and hyperlink, we only apply them to protect the email address, as discussed in Section 7. In particular, we transform the email address in each personal profile in the synthetic dataset into the corresponding format of each defense. Regarding text-to-image, we use an open-source tool [5] to convert personal information into an image and embed it into the corresponding profile.
For prompt injection defense, we inject the prompt shown in Table 18 in the Appendix to all personal profiles in the synthetic dataset and Figure 7 in the Appendix presents how we make the injection. By default, we use “combination of context ignoring and injected data (CI+ID)”.
8.2 Defense Results
Prompt injection outperforms existing defenses :  Table 8 shows the attack results to extract the email address when different defenses (excluding text-to-image) are adopted. For symbol replacement, we consider three variants: replacing “@” only, replacing “.” only, and replacing both. We observe that prompt injection is the most effective defense, as the accuracy of the email address extracted by the attacker drops drastically when prompt injection is applied. This indicates that the injected prompt can successfully deceive the LLMs. By contrast, other defenses only have limited effectiveness, as we only observe a slight decrease when these defenses are used. In addition, Table 17 in the Appendix shows the experimental results of the text-to-image defense when multi-modal LLMs (i.e., GPT-4V and Gemini-pro-vision) are used. The defense only decreases the attack effectiveness slightly, which indicates that text-to-image is not effective in preventing extraction based on multi-modal LLMs. Overall, the experimental results demonstrate that prompt injection outperforms other countermeasures when defending against LLM-based email address extraction from personal profiles.
Table 9 shows the results to extract the email address of different extraction methods (i.e., traditional or LLM-based) under different defenses. We have the following observations. First, the LLM-based method achieves better or at least comparable results compared to traditional methods when prompt injection is not used. Second, prompt injection can effectively defend against LLM-based methods, reducing the attack to traditional ones. In particular, when prompt injection is adopted, the performance of the LLM-based method decreases to 0, while the performance of traditional methods is close to those without defense. We notice that there is a 1% difference for the accuracy of regular expression under no-defense and prompt injection. This is because when the ground-truth label is “no email”, regular expression extracts correctly when no defense is used, but extracts the injected prompt when prompt injection is adopted, which reduces the accuracy slightly. Moreover, when prompt injection is used along with existing defense (i.e., keyword replacement), it successfully defends against LLM-based and traditional PIE methods, i.e., the performance of all methods becomes zeros.
We find that the Rouge1 score (or BERT score) for the work and educational experience extracted by the attacker are not close to zero, but they remain at around 0.2 (or 0.25). We suspect this is due to the design of our injected prompt. As shown in Table 18, the injected prompt aims to mislead the LLM to answer with “imaginary company” (or “imaginary school”) when the LLM is originally guided by the attacker to extract the work experience (or education experience). Thus, even if the injected prompt successfully deceives the LLM and makes it generate “imaginary company” (or ‘imaginary school”), the keyword “company” (or “school”) is very likely to have some overlaps with the words in the ground truth labels, resulting in non-zero Rouge1 scores (or BERT scores).
@@ 8.2 Defense Results
Data prompt isolation (quotes) w/o defense 50% / 0.50 100% / 1.00 91% / 0.70 100% / 1.00 61% / 0.46 74% / 0.55 91% / 0.88 50% / 0.39
@@
Data prompt isolation (xml tags) w/o defense 99% / 0.99 99% / 0.99 78% / 0.76 100% / 1.00 72% / 0.58 78% / 0.65 89% / 0.86 55% / 0.44
@@
Data prompt isolation (random sequence) w/o defense 99% / 0.99 99% / 0.99 77% / 0.75 100% / 1.00 67% / 0.50 76% / 0.58 89% / 0.86 56% / 0.46
@@
Instructional prevention w/o defense 99% / 0.99 96% / 0.96 78% / 0.76 100% / 1.00 65% / 0.47 74% / 0.55 89% / 0.85 53% / 0.42
@@
Sandwich prevention w/o defense 99% / 0.99 99% / 0.99 78% / 0.77 100% / 1.00 65% / 0.49 75% / 0.57 89% / 0.87 54% / 0.43
@@
Impact of the prompt used in prompt injection:  Table 21 in the Appendix shows the impact of prompt injection strategies. The prompts used in our experiment when only one strategy is adopted are shown in Table 18 in the Appendix. We observe that LLM-based PIE is still very effective when only one strategy is applied. By contrast, the accuracy (or Rouge1/BERT score) is small when both strategies are used. This highlights the importance of both the context ignoring strategy and the injected data in designing the injected prompt to effectively defend against LLM-based PIE.
@@
Prompt injection is effective in defending against adaptive attacks:  Existing works propose various ways to defend against prompt injection attacks [7, 48, 54, 35, 31]. Their main goal is to prevent the LLMs from being misled by the injected prompt via either purifying the entire prompt or separating the instruction and the document/personal profile. In our scenario, we utilize them as adaptive attacks to break the defensive prompt injection used by the personal profile owners. In other words, these adaptive attacks aim to make the LLMs accurately extract the personal information from the profile, instead of being misguided by the injected prompt. We utilize paraphrasing, retokenization, data prompt isolation, instructional prevention, and sandwich prevention as adaptive attacks. A summary of adaptive attacks and examples are in Table 22 in the Appendix. In our evaluation, we consider our attack framework as the baseline attack, and integrate these adaptive attacks into our attack framework one at a time. In addition, we do not consider prompt injection detection methods as adaptive attacks in our evaluation, since even with the knowledge that a document contains injected prompts, the attacker still cannot distinguish the genuine and injected contents, and thus fails to extract accurate personal information.
@@
Table 11 shows the attack results for these adaptive attacks. We have two observations. First, some adaptive attacks, e.g., retokenization, achieve a higher attack accuracy (or Rouge1/BERT scores) compared to baselines. However, the increase from these adaptive attacks is very limited. For instance, when extracting the email address, the retokenization only achieves an accuracy of 0.22. Second, we observe that some adaptive attacks sacrifice the utility. In particular, when the prompt injection is not applied, the attack results for some adaptive attacks are lower than the baseline attack. For example, when extracting the email address, the retokenization only has an accuracy of 0.48, which is much lower than the value of the baseline attack (i.e., 0.93). Based on two observations, we could conclude that these adaptive attacks are insufficient in bypassing the prompt injection defense.
9 Discussion
Computational cost:  Table 16 in the Appendix compares the average runtime of various PIE methods for extracting a category of personal information from a single profile in our synthetic dataset. LLM-based methods, such as GPT-4, incur higher computational costs compared to traditional methods. We suspect that the computational overhead is primarily due to a large number of parameters of LLMs. While traditional methods are more efficient, they are less effective than LLM-based methods as shown in our results. Moreover, we note that it takes around 1 second for the LLM-based method, which is acceptable in practice.
Financial cost:  We also evaluate the financial costs of LLM-based PIE methods. In particular, extracting a single category of personal information from a personal profile (or image) using the Microsoft Azure GPT-4 API incurs a query cost of \$0.009 (or \$0.013). To reduce financial costs, an attacker can use open-source LLMs, though they can be less effective in extracting personal information as shown in our experimental results. In other words, there can be a trade-off between the financial cost and attack effectiveness.
Comparision with crowdsourcing:  Crowdsourcing, such as Amazon Mechanical Turk (MTurk) that leverages human annotators to perform labeling tasks, is another traditional way of information extraction. However, they can have the following limitations: 1) potential inconsistencies in annotator performance, 2) higher economic cost than LLM-based PIE, and 3) unreliable and low-quality results [47]. For instance, in terms of financial cost, the average wage for a worker on MTurk is approximately \$2 per hour [28]. Based on our manual annotation of 100 profiles from real-world datasets, where each profile includes 7 categories of personal information, we estimate a total annotation time of 4 hours. This results in a cost of \$0.011 per profile per category. In comparison, GPT-4 requires only \$0.009 to extract one category of personal information per profile, demonstrating that LLMs are a more cost-efficient alternative to crowdsourcing.
Memorization test:  A critical concern in evaluating LLM-based methods is their potential reliance on memorized training data, which could artificially inflate performance when the datasets used for evaluation overlap with the model’s pretraining corpus. Following [53], we conducted a memorization test for the Synthetic, Celebrity, Physician, Professor, and Court datasets using GPT-4. We calculated string similarity ratios between the dataset samples and GPT-4’s outputs. For each dataset, we measured the proportion of samples with similarity ratios below 0.6, which indicates that the model is not likely to have memorized the data. The results were 98%, 100%, 96%, 100%, and 100% for the five datasets, respectively, demonstrating that GPT-4 is unlikely to have memorized them.
Prompt injection usability:  Prompt injection as a defense may influence the readability and accessibility of webpage content. For instance, making injected prompts invisible to human readers (e.g., by matching the font color to the background) may disrupt accessibility tools like screen readers. These tools may read invisible prompts aloud, thereby creating barriers for users with disabilities. We believe it is an interesting future work to develop prompt injection techniques to minimize interference with assistive technologies while maintaining robust protection against LLM-based PIE attacks.
Adaptive attacks:  While our results show that prompt injection is effective in misleading LLM-based PIE, even against adaptive attacks (as demonstrated in Table 11), there are more potential adaptive attacks to consider. For instance, instead of parsing text, attackers could capture screenshots of personal profiles. In such cases, the invisible injected prompts would not be included in the screenshots, allowing the LLM to potentially bypass the defense. One possible countermeasure could involve image-based prompt injection, where the website developers incorporate tailored visual perturbations that are imperceptible to human users but crafted to mislead multimodal LLMs capable of processing images. This approach could ensure that even if attackers rely on screenshots, the personal information extraction can be misled. Further research can focus on exploring the feasibility and effectiveness of such defenses against multimodal LLMs.